Combined Communications is a new firm in a rapidly growing industry. The company is planning on increasing its annual dividend by 21 percent a year for the next 4 years and then decreasing the growth rate to 6 percent per year. The company just paid its annual dividend in the amount of $1.30 per share. What is the current value of one share of this stock if the required rate of return is 9.50 percent?

$70.61

$60.88

$65.42

$64.46

$56.60
Answer #1

Answer ; $ 65.42

a. Computation of present value of expected dividends:

Year Expected Dividend PV factor at 9.5 % PV of Dividends
1 $ 1.30 * ( 1.21 ) 1 0.913 $ 1.44
2 $ 1.30 * ( 1.21 ) 2 0.834 1.59
3 $ 1.30 * ( 1.21 ) 3 0.762 1.75
4 $ 1.30 * ( 1.21) 4 0.696 1.94
Total $ 6.72

b. Computation of present value at the end of Year 4 :

PV = $ 1.30 * ( 1.21 ) 4 * ( 1.06 ) / ( 0.095 - 0.06 ) = $ 2.9539 / 0.035 = $ 84.40

c. Computation of present value today at 9.5 % discount rate:

P0 = $ 84.40 / ( 1.095 ) 4 = $ 84.40 x 0.696 = $ 58.74

d. Present value of the stock = $ 6.72 + $ 58.74 = $ 65.46